OpenStudy (anonymous):

Y+6=-3/4(x+8) Write in slope intercept form?

OpenStudy (anonymous):

\[y+6=\frac{ -3 }{ 4 }(x+8)\] \[y+6=\frac{ -3 }{ 4 }x-6\] \[y=\frac{ -3 }{ 4 }x-6-6\] \[y=\frac{ -3 }{ 4}x -12\]